Project

General

Profile

Bug #1788

Muon Beamline Step IV Prep list

Added by Nebrensky, Henry over 8 years ago. Updated about 8 years ago.

Status:
Closed
Priority:
Normal
Category:
Beam Line
Start date:
28 October 2015
Due date:
16 February 2016
% Done:

100%

Estimated time:

Description

Issues that need sorting out before Step IV


Related issues

Follows Operations - Bug #1708: Muon Beamline Step IV prep listClosedNebrensky, Henry13 July 201508 September 2015

Actions
Precedes Operations - Feature #1854: Muon Beamline Step IV prep listClosedNebrensky, Henry06 June 201628 June 2016

Actions
#1

Updated by Nebrensky, Henry over 8 years ago

Stuff that needs sorting out by 23rd February 2016, c.f. January maintenance list
Important stuff in bold

Target

  • Light Loss (TGT) - stabilised in mild saturation, now falling again
  • Trace fibre connections - January shutdown awaiting MICE Note
  • Restart loopback measurements (TGT)
  • Protocol for synch access for swapping fibres at patch panel (TGT)
  • S3 and T2 - assembly of next spare (TGT) - pinch point is getting optics block off T1
  • Clean-up of R78 target control crate (DL)
  • Shifter "Target Drive" panel (CAM) - needs testing in MDR6
  • Target Alarm handler #1613 update 22 (CAM)
  • ISIS data items #1669 (CAM)
  • Telephone link #977
  • Activation Study (activation seen: limit of 2 V at MS/64 stands)
  • Beambump tune
  • Updated target manual (EO/JL)
  • Check Target Controller for USB Instability/Over-current

Conventional Magnets

  • Second chiller pump or hard limit in controls (VC175 )
    • review existing water flows (MBC)
    • need limit in control software? (CAM)
  • Q4 oscillation below 10A and anomalous ramp rate (DL)
  • Q1 MPS start-up recalcitrance (DL)
  • Q1 alarm on voltage at zero amps (CAM)
  • Q4 Earth Leakage (MBC) - D2,Q4,Q5 flushed through
  • Q6 leak leave until time
  • Leak sensors and drip trays
  • "All to 0A" controls issue
  • Water, temperatures and MPS water flows in ALH (CAM)
  • Q5 and Q6 temperature sensors swapped (CAM)
  • POMPOM readout into EPICS (CAM)
  • Update magnet and POMPOM user guides (MBC)
  • Annual POMPOM radiation damage test (MBC)
    • negative
    • positive

Decay Solenoid

  • Overheating investigation (DK/DL)
  • Replace PSU (DK/DL)
    • Calibrate current reading and update IOC software (DL)
  • 120A trip investigation (DK/DL)
  • Replace final battery cell (DL -confirmed still outstanding )
  • Diodes in BBU circuit (DL)
  • Boosted air (MBC)
    • Trench sensor (awaiting read-out)
    • Compressor sensor
    • Backup compressor
  • Linde PC install: awaiting validation and IOC (DK) - licence issues
    • MLCR monitor and KVM implications (HN)
  • Vac pump tables
  • ALH/archiver corrections #1710 (CAM)
  • Off-site EPICS Gateway/Archiver access #1460 / #1740
  • Power control fix / "All to 0A" controls issue (CAM)
  • Expert-only mode: password (CAM)
  • Magnet-On status lamp (DL) - needs synch access

Proton Absorber

  • Combine survey documents (MBC)
  • 15mm slab

Beamstop

  • MICE-ISIS Safety mtg permission (MBC / Paul Wright)
  • Translate Survey document (MBC)

General

  • BLOC training and experience (MBC)
  • Shared services/environment (water, air, computing/networking, UPS ( #1213 ) in Alarm Handler and in Archiver (CAM)

Beyond the Muon Beamline

#2

Updated by Nebrensky, Henry about 8 years ago

  • Status changed from Open to Closed
  • % Done changed from 0 to 100

Close and hand over to #1854

Also available in: Atom PDF